What is the maximum value of the transformed function: g(x)=−3cos[2(x−60°)]+3?
-6
3
6
-3

Found 2 solutions by ikleyn, greenestamps:
Answer by ikleyn(53751)

.

The maximum of the function is reached when cosine is equal to -1,


and this maximum is  (-3)*(-1) + 3 = 3 + 3 = 6.    ANSWER

Solved.


Answer by greenestamps(13327)



A different approach....

The "+3" at the end of the function tells us the midline of the graph of the function is y=3.

The factor "-3" on the cosine function tells us the function value varies 3 units either side of the midline; that makes the maximum and minimum function values 3+3=6 and 3-3=0

ANSWER: maximum value: 6